# Decision making in stochastic extensive form I: Stochastic decision forests

### v1

@@ A.4 Section 4
Regarding AP.C1, let $w=(\omega,f)\in c$. There is component-wise decreasing $f^{\prime}\in\mathbb{A}^{\mathbb{T}}$ with $f^{\prime}|_{[0,t)_{\mathbb{T}}}=f|_{[0,t)_{\mathbb{T}}}$ and $p^{i}\circ f^{\prime}(t)\neq g(\omega)$, because $p^{i}\circ f|_{[0,t)_{\mathbb{T}}}=1_{[0,t)_{\mathbb{T}}}$. Then $w^{\prime}=(\omega,f^{\prime})\in W$, hence $w^{\prime}\in x_{t}(w)$, but $w^{\prime}\notin c$.
@@
Regarding AP.C2, let $f\in\mathbb{A}^{\mathbb{T}}$ with $f|_{[0,t)_{\mathbb{T}}}\in A_{<t}$ and $\omega\in D_{t,f}$. There is component-wise decreasing $f^{\prime}\in\mathbb{A}^{\mathbb{T}}$ with $f^{\prime}|_{[0,t)_{\mathbb{T}}}=f|_{[0,t)_{\mathbb{T}}}$ and $p^{i}\circ f^{\prime}(t)=g(\omega)$, by assumption on $A_{<t}$. As $(\omega,f^{\prime})\in W$, we infer
@@
Regarding assumption AP.C3, let ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ be such that, with $t=\mathfrak{t}({\bm{\mathsf{x}}})$, there are $D\in\mathscr{A}$, $g\colon D\to\mathbb{A}^{i}=\{0,1\}$ and $A_{<t}\subseteq\mathbb{A}^{[0,t)_{\mathbb{T}}}$ satisfying $c(A_{<t},i,g)\in\mathscr{C}_{t}$ and making $c(A_{<t},i,g)$ available at ${\bm{\mathsf{x}}}$. By lemma 4.38, there is $(\omega_{0},f_{0})\in c(A_{<t},i,g)$ such that $\omega_{0}\in D_{t,f_{0}}=D_{\bm{\mathsf{x}}}$ and ${\bm{\mathsf{x}}}={\bm{\mathsf{x}}}_{t}(f_{0})$. As $c(A_{<t},i,g)\in\mathscr{C}_{t}$, assumption AP.C1 implies, that $p^{i}\circ f_{0}|_{[0,t)_{\mathbb{T}}}$ is constant with value $1$. Indeed, there is $f_{1}\in\mathbb{A}^{\mathbb{T}}$ with $(\omega,f_{1})\in x_{t}(\omega,f_{0})\setminus c(A_{<t},i,g)$. In particular, $f_{1}$ is component-wise decreasing and takes the same values on $[0,t)_{\mathbb{T}}$ as $f_{0}$. If $p^{i}\circ f_{0}$ took the value 00 at some point in $[0,t)_{\mathbb{T}}$, the monotonicity would imply $p^{i}\circ f_{1}(t)=0=p^{i}\circ f_{0}(t)=g(\omega)$, whence $(\omega,f_{1})\in c(A_{<t},i,g)$, in contradiction to the choice of $f_{1}$.
@@
Let $A^{\prime}_{<t}$ be the set of component-wise decreasing $f_{t}\in\mathbb{A}^{[0,t)_{\mathbb{T}}}$ such that $p^{i}\circ f_{t}$ is constant with value $1$. As $A_{<t}\subseteq A^{\prime}_{<t}$, property AP.C3(1) is clearly satisfied. Further, let $\mathscr{G}(\mathbb{A}^{i})=\mathscr{B}(\mathbb{A}^{i})\setminus\{\mathbb{A}^{i}\}$ which is obviously an intersection-stable generator of $\mathscr{B}(\mathbb{A}^{i})$. As $\mathbb{A}^{i}=\{0,1\}$, $\mathscr{G}(\mathbb{A}^{i})=\{\{0\},\{1\},\emptyset\}$. Let $G\in\mathscr{G}(\mathbb{A}^{i})$. If $G=\emptyset$, then $c(A^{\prime}_{<t},A_{t}^{i,G})=\emptyset$. Else, that is, if $G$ is a singleton, then $c(A^{\prime}_{<t},A_{t}^{i,G})=c(A^{\prime}_{<t},i,g_{G})$ for the constant map $g_{G}$ with value given by the unique element of $G$. We have shown just beforehand that $c(A^{\prime}_{<t},i,g_{G})\in\mathscr{C}_{t}$. To complete the proof of the fact that $c(A^{\prime}_{<t},A_{t}^{i,G})\in\mathscr{C}_{\bm{\mathsf{x}}}$, it thus remains to show axiom (4) in the definition of $\mathscr{C}_{\bm{\mathsf{x}}}$, because (3) has just been proven and (1) and (2) are evident by construction. For this, let $\omega^{\prime}\in D_{\bm{\mathsf{x}}}$. As $p^{i}\circ f_{0}$ only takes the value $1$ on $[0,t)_{\mathbb{T}}$, there is component-wise decreasing $f^{\prime}\colon\mathbb{T}\to\mathbb{A}$ such that $f^{\prime}|_{[0,t)_{\mathbb{T}}}=f_{0}|_{[0,t)_{\mathbb{T}}}$ and $p^{i}\circ f^{\prime}(t)=g_{G}(\omega^{\prime})$. Hence, $(\omega^{\prime},f^{\prime})\in{\bm{\mathsf{x}}}_{t}(f_{0})(\omega^{\prime})\cap c(\tilde{A}^{\prime}_{<t},i,g_{G})={\bm{\mathsf{x}}}(\omega^{\prime})\cap c(A^{\prime}_{<t},A_{t}^{i,G})$. This completes the proofs for the timing game example.
@@
(Ad up-and-out option exercise example):  Let $t\in\mathbb{R}_{+}$, $A_{<t}=\{1\}^{\mathbb{R}_{+}}$ and $D$ be the set of $\omega\in\Omega$ such that $\max_{u\in[0,t]}P_{u}(\omega)<2$. We suppose that $D\neq\emptyset$. Let $g\colon D\to\{0,1\}$ be a map and let $c=c(A_{<t},i,g)$. We are going to prove that $c\in\mathscr{C}_{t}$.
@@
Regarding assumption AP.C3, let ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ be such that, with $t=\mathfrak{t}({\bm{\mathsf{x}}})$, there are $\tilde{D}\in\mathscr{A}$, $g\colon\tilde{D}\to\mathbb{A}^{i}=\{0,1\}$ and $A_{<t}\subseteq\mathbb{A}^{[0,t)}$ satisfying $c(A_{<t},i,g)\in\mathscr{C}_{t}$ and making $c(A_{<t},i,g)$ available at ${\bm{\mathsf{x}}}$. By lemma 4.38, there is $(\omega_{0},f_{0})\in c(A_{<t},i,g)$ such that ${\bm{\mathsf{x}}}={\bm{\mathsf{x}}}_{t}(f_{0})$ and $\omega_{0}\in D_{\bm{\mathsf{x}}}$. As $c(A_{<t},i,g)\in\mathscr{C}_{t}$, assumption AP.C1 implies, that $f_{0}|_{[0,t)}$ is constant with value $1$. Indeed, there is $f_{1}\in\mathbb{A}^{\mathbb{T}}$ with $(\omega,f_{1})\in x_{t}(\omega,f_{0})\setminus c(A_{<t},i,g)$. In particular, $f_{1}$ is component-wise decreasing and takes the same values on $[0,t)_{\mathbb{T}}$ as $f_{0}$. If $f_{0}$ took the value 00 at some point in $[0,t)$, the monotonicity would imply $f_{1}(t)=0=f_{0}(t)=g(\omega)$, whence $(\omega,f_{1})\in c(A_{<t},i,g)$, in contradiction to the choice of $f_{1}$.
Let $A^{\prime}_{<t}=\{1\}^{[0,t)}$. As $A_{<t}\subseteq A^{\prime}_{<t}$, property AP.C3(1) is clearly satisfied. Further, let $\mathscr{G}(\{0,1\})=\mathscr{B}(\{0,1\})\setminus\{\{0,1\}\}$ which is obviously an intersection-stable generator of ($\mathscr{B}(\{0,1\})$. Note that $\mathscr{G}(\{0,1\})=\{\{0\},\{1\},\emptyset\}$. Let $G\in\mathscr{G}(\{0,1\})$. If $G=\emptyset$, then $c(A^{\prime}_{<t},A_{t}^{i,G})=\emptyset$. Else, that is, if $G$ is a singleton, then $c(A^{\prime}_{<t},A_{t}^{i,G})=c(A^{\prime}_{<t},i,g_{G})$ for the constant map $g_{G}$ with value given by the unique element of $G$.
We have shown just above that $c(A^{\prime}_{<t},i,g_{G})\in\mathscr{C}_{t}$ because $A^{\prime}_{<t}\subseteq\mathbb{A}^{[0,t)}$ is a non-empty set of decreasing paths $f_{t}$ such that $f_{t}=1_{[0,t)}$. Hence, $c(A^{\prime}_{<t},i,g_{G})\in\mathscr{C}_{t}$. To complete the proof of the fact that $c(A^{\prime}_{<t},i,g_{G})\in\mathscr{C}_{\bm{\mathsf{x}}}$, it remains to show axiom (4) in the definition of $\mathscr{C}_{\bm{\mathsf{x}}}$, because (3) has just been proven and (1) and (2) are evident by construction. For this, let $\omega^{\prime}\in D_{\bm{\mathsf{x}}}=D_{t,f_{0}}$. In particular, $\max_{u\in[0,t]}P_{u}(\omega^{\prime})<2$. By continuity of $P$, there is $\varepsilon>0$ such that $\max_{u\in[0,t+\varepsilon]}P_{u}(\omega^{\prime})<2$. Then, regardless of the value of $g(\omega)$, and as $f_{0}$ only takes the value $1$ on $[0,t)$, there is component-wise decreasing $f^{\prime}\colon{\mathbb{R}_{+}}\to\mathbb{A}$ such that $f^{\prime}|_{[0,t)}=f_{0}|_{[0,t)}$, $f^{\prime}(t)=g_{G}(\omega^{\prime})$, and $f^{\prime}(t+\varepsilon)=0$. Hence, $(\omega^{\prime},f^{\prime})\in{\bm{\mathsf{x}}}(\omega^{\prime})\cap c(\tilde{A}^{\prime}_{<t},i,g_{G})$. This completes the proofs for the up-and-out option exercise problem example. ∎

### v2

@@ A.4 Section 4
Regarding AP.C1, let $w=(\omega,f)\in c$. There is componentwise decreasing $f^{\prime}\in\mathbb{A}^{\mathbb{T}}$ with $f^{\prime}|_{[0,t)_{\mathbb{T}}}=f|_{[0,t)_{\mathbb{T}}}$ and $p^{i}\circ f^{\prime}(t)\neq g(\omega)$, because $p^{i}\circ f|_{[0,t)_{\mathbb{T}}}=1_{[0,t)_{\mathbb{T}}}$. Then $w^{\prime}=(\omega,f^{\prime})\in W$, hence $w^{\prime}\in x_{t}(w)$, but $w^{\prime}\notin c$.
@@
Regarding AP.C2, let $f\in\mathbb{A}^{\mathbb{T}}$ with $f|_{[0,t)_{\mathbb{T}}}\in A_{<t}$ and $\omega\in D_{t,f}$. There is componentwise decreasing $f^{\prime}\in\mathbb{A}^{\mathbb{T}}$ with $f^{\prime}|_{[0,t)_{\mathbb{T}}}=f|_{[0,t)_{\mathbb{T}}}$ and $p^{i}\circ f^{\prime}(t)=g(\omega)$, by assumption on $A_{<t}$. As $(\omega,f^{\prime})\in W$, we infer
@@
Regarding Assumption AP.C3, suppose that for all ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ with $\mathfrak{t}({\bm{\mathsf{x}}})=t$, $\mathscr{C}^{i}_{\bm{\mathsf{x}}}$ contains all $c(A_{<t},A_{t})$ ranging over all $A_{t}$ satisfying ($\mathscr{C}_{\bm{\mathsf{x}}}^{i}$.$k$), $k=2,3,4$. Let ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ be such that $c(A_{<t},i,g)$ is available at ${\bm{\mathsf{x}}}$. By Lemma 4.41, there is $(\omega_{0},f_{0})\in c(A_{<t},i,g)$ such that $\omega_{0}\in D_{t,f_{0}}=D_{\bm{\mathsf{x}}}$ and ${\bm{\mathsf{x}}}={\bm{\mathsf{x}}}_{t}(f_{0})$. As $c(A_{<t},i,g)\in\mathscr{C}_{t}$, Assumption AP.C1 implies, that $p^{i}\circ f_{0}|_{[0,t)_{\mathbb{T}}}$ is constant with value $1$. Indeed, there is $f_{1}\in\mathbb{A}^{\mathbb{T}}$ with $(\omega,f_{1})\in x_{t}(\omega,f_{0})\setminus c(A_{<t},i,g)$. In particular, $f_{1}$ is componentwise decreasing and takes the same values on $[0,t)_{\mathbb{T}}$ as $f_{0}$. If $p^{i}\circ f_{0}$ took the value 00 at some point in $[0,t)_{\mathbb{T}}$, the monotonicity would imply $p^{i}\circ f_{1}(t)=0=p^{i}\circ f_{0}(t)=g(\omega)$, whence $(\omega,f_{1})\in c(A_{<t},i,g)$, in contradiction to the choice of $f_{1}$.
@@
Further, let $\mathscr{G}(\mathbb{A}^{i})=\mathscr{B}(\mathbb{A}^{i})\setminus\{\mathbb{A}^{i},\emptyset\}$ which is obviously a generator of $\mathscr{B}(\mathbb{A}^{i})$ stable under non-trivial intersections. As $\mathbb{A}^{i}=\{0,1\}$, $\mathscr{G}(\mathbb{A}^{i})=\{\{0\},\{1\}\}$. Let $G\in\mathscr{G}(\mathbb{A}^{i})$. Thus, $G$ is a singleton and $c(A_{<t},A_{t}^{i,G})=c(A_{<t},i,g_{G})$ for the constant map $g_{G}$ with value given by the unique element of $G$. We have shown just beforehand that $c(A_{<t},i,g_{G})\in\mathscr{C}_{t}$. To complete the proof of the fact that $c(A_{<t},A_{t}^{i,G})\in\mathscr{C}_{\bm{\mathsf{x}}}^{i}$, in view of our additional assumption on $\mathscr{C}^{i}_{\bm{\mathsf{x}}}$, it thus remains to show Axiom ($\mathscr{C}_{\bm{\mathsf{x}}}^{i}$.4) in the definition of $\mathscr{C}_{\bm{\mathsf{x}}}$, because ($\mathscr{C}_{\bm{\mathsf{x}}}^{i}$.3) has just been proven and ($\mathscr{C}_{\bm{\mathsf{x}}}^{i}$.2) is evident by construction. For this, let $\omega^{\prime}\in D_{\bm{\mathsf{x}}}$. As $p^{i}\circ f_{0}$ only takes the value $1$ on $[0,t)_{\mathbb{T}}$, there is componentwise decreasing $f^{\prime}\colon\mathbb{T}\to\mathbb{A}$ such that $f^{\prime}|_{[0,t)_{\mathbb{T}}}=f_{0}|_{[0,t)_{\mathbb{T}}}$ and $p^{i}\circ f^{\prime}(t)=g_{G}(\omega^{\prime})$. Hence, $(\omega^{\prime},f^{\prime})\in{\bm{\mathsf{x}}}_{t}(f_{0})(\omega^{\prime})\cap c(A_{<t},i,g_{G})={\bm{\mathsf{x}}}(\omega^{\prime})\cap c(A_{<t},A_{t}^{i,G})$. This completes the proofs for the timing game example.
@@
(Ad up-and-out option exercise example):  Let $t\in\mathbb{R}_{+}$, $A_{<t}=\{1\}^{[0,t)}$ and $D$ be the set of $\omega\in\Omega$ such that $\max_{u\in[0,t]}P_{u}(\omega)<2$. We suppose that $D\neq\emptyset$. Let $g\colon D\to\{0,1\}$ be a map and let $c=c(A_{<t},i,g)$. We are first going to prove that $c\in\mathscr{C}_{t}$.
@@
Regarding Assumption AP.C3, suppose that for all ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ with $\mathfrak{t}({\bm{\mathsf{x}}})=t$, $\mathscr{C}^{i}_{\bm{\mathsf{x}}}$ contains all $c(A_{<t},A_{t})$ ranging over all $A_{t}$ satisfying ($\mathscr{C}_{\bm{\mathsf{x}}}^{i}$.$k$), $k=2,3,4$. Let ${\bm{\mathsf{x}}}\in\tilde{\bm{\mathsf{X}}}^{i}$ be such that $c(A_{<t},i,g)$ is available at ${\bm{\mathsf{x}}}$. By Lemma 4.41, there is $(\omega_{0},f_{0})\in c(A_{<t},i,g)$ such that ${\bm{\mathsf{x}}}={\bm{\mathsf{x}}}_{t}(f_{0})$ and $\omega_{0}\in D_{\bm{\mathsf{x}}}$. As $c(A_{<t},i,g)\in\mathscr{C}_{t}$, Assumption AP.C1 implies, that $f_{0}|_{[0,t)}$ is constant with value $1$. Indeed, there is $f_{1}\in\mathbb{A}^{\mathbb{T}}$ with $(\omega,f_{1})\in x_{t}(\omega,f_{0})\setminus c(A_{<t},i,g)$. In particular, $f_{1}$ is componentwise decreasing and takes the same values on $[0,t)_{\mathbb{T}}$ as $f_{0}$. If $f_{0}$ took the value 00 at some point in $[0,t)$, the monotonicity would imply $f_{1}(t)=0=f_{0}(t)=g(\omega)$, whence $(\omega,f_{1})\in c(A_{<t},i,g)$, in contradiction to the choice of $f_{1}$.
Let $\mathscr{G}(\{0,1\})=\mathscr{B}(\{0,1\})\setminus\{\{0,1\},\emptyset\}$ which is obviously a generator of ($\mathscr{B}(\{0,1\})$ stable under non-trivial intersections. Note that $\mathscr{G}(\{0,1\})=\{\{0\},\{1\}\}$. Let $G\in\mathscr{G}(\{0,1\})$. Thus, $G$ is a singleton and $c(A_{<t},A_{t}^{i,G})=c(A_{<t},i,g_{G})$ for the constant map $g_{G}$ with value given by the unique element of $G$.
We have shown just above that $c(A_{<t},i,g_{G})\in\mathscr{C}_{t}$. To complete the proof of the fact that $c(A_{<t},i,g_{G})\in\mathscr{C}_{\bm{\mathsf{x}}}^{i}$, in view of our additional assumption on $\mathscr{C}^{i}_{\bm{\mathsf{x}}}$, it remains to show Axiom ($\mathscr{C}_{\bm{\mathsf{x}}}^{i}$.4), because ($\mathscr{C}_{\bm{\mathsf{x}}}^{i}$.3) has just been proven and ($\mathscr{C}_{\bm{\mathsf{x}}}^{i}$.2) is evident by construction. For this, let $\omega^{\prime}\in D_{\bm{\mathsf{x}}}=D_{t,f_{0}}$. In particular, $\max_{u\in[0,t]}P_{u}(\omega^{\prime})<2$. By continuity of $P$, there is $\varepsilon>0$ such that $\max_{u\in[0,t+\varepsilon]}P_{u}(\omega^{\prime})<2$. Then, regardless of the value of $g(\omega)$, and as $f_{0}$ only takes the value $1$ on $[0,t)$, there is componentwise decreasing $f^{\prime}\colon{\mathbb{R}_{+}}\to\mathbb{A}$ such that $f^{\prime}|_{[0,t)}=f_{0}|_{[0,t)}$, $f^{\prime}(t)=g_{G}(\omega^{\prime})$, and $f^{\prime}(t+\varepsilon)=0$. Hence, $(\omega^{\prime},f^{\prime})\in{\bm{\mathsf{x}}}(\omega^{\prime})\cap c(A_{<t},i,g_{G})$. This completes the proofs for the up-and-out option exercise problem example. ∎